# Attention as an Adaptive Filter

### v2

@@ 7 Discussion and Conclusion
This work explores a connection between classical control theory and attention mechanisms by framing attention through the lens of linear dynamical systems and adaptive filtering. We show that attention arises naturally as the solution to a maximum likelihood estimation problem under a linear stochastic differential equation with structured system matrices and noise covariances, and discrete observations. Unlike approaches that treat attention as a static kernel or a matrix factorization, we view attention as a parallelized form of adaptive Kalman Filtering: each key-value pair is interpreted as a noisy observation of a latent state, and a precision-weighted Bayesian update aggregates evidence across all observations to refine the state estimate in a parallel manner. To the best of our knowledge, this view of attention as the MLE for a structured linear SDE has not appeared in the literature on attention mechanisms.
Imposing an isotropic-plus-skew-symmetric state matrix and isotropic noise, our model achieves the same asymptotic time and memory complexity as standard attention. This result in a form of complex-valued rotary positional encoding equivalent to that found in xPos and RetNet, along with a query-key precision kernel that quantifies an uncertainty prior under the learned dynamics model.
@@
We also introduce a Radial–Tangential SDE model, whose isotropic-plus-skew-symmetric state matrix and anisotropic noise in the token frame define the most expressive SDE structure that remains analytically tractable. This formulation leads directly to a maximum likelihood estimator on the hypersphere, which provides a novel, principled foundation for understanding norm-preserving attention mechanisms, shedding new light on the architectural properties of the Transformer.
@@
Several promising directions remain for future work. A natural extension is to allow the state matrix and noise to depend on the state. Although a fully time-varying dynamics model would break the linear time-invariant (LTI) structure and would no longer allow a closed-form solution to the Differential Lyapunov Equation (DLE), this issue can be avoided by fixing the dynamics within each context window. In this way, the parameters may adapt across windows while remaining constant within a window, enabling content adaptivity without sacrificing computational efficiency.
@@
The closed-form expression for the propagated precision matrix, which assumes simultaneous diagonalizability, can be computed in $\mathcal{O}(md)$ time, whereas full integration and inversion in the general case requires $\mathcal{O}(md^{3})$ time. Since our algorithm already involves $\mathcal{O}(md^{2})$ matrix multiplications, moderately less efficient integration methods introduce a limited additional overhead. In particular, if we use a diagonalizable-plus-rank-1 state matrix $\mathbf{A}$, covariance propagation can be computed efficiently by expressing the matrix exponential and integrals as a diagonal term plus a rank-1 outer-product correction, reducing complexity to $\mathcal{O}(md^{2})$, though losing a fully closed form. We could also consider a generalization using the Jordan canonical form. Alternatively, differentiable numerical solvers for the differential Lyapunov equation, or unrolled iterations of specialized low-rank methods such as Krylov subspace techniques or ADI iterations could potentially be integrated directly into a neural network layer.
Finally, the duality between attention weights and the Kalman gain suggests a path toward dynamics-aware neural control policies, where the separation principle can be leveraged to achieve more sample-efficient policy evaluation and state estimation in noisy or partially observed reinforcement learning settings.
@@ Appendix A Background
A.1 Propagation of Uncertainty
@@ A.3 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\sigma_{V}^{2}(\Delta t_{ij})\mathbf{I},\quad\text{where}\quad\sigma_{V}^{2}(\Delta t_{ij}):=\text{sign}(\Delta t_{ij})\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}+e^{-2\alpha\Delta t_{ij}}\eta^{2}$
@@ A.4 Variational Formulation of Robust Estimation
Dividing numerator and denominator by $\beta$ and setting the numerator to $1$ (since it is canceled by the normalization), there is only one independent constant, $\nu=2\beta$, which controls the residual scaling:
@@
The term $\mathbf{r}^{\top}\mathbf{V}^{-1}\mathbf{r}$ is known as the Mahalanobis distance, which quantifies the difference between a vector $\mathbf{r}$ and a distribution with covariance $\mathbf{V}$. Here, $\mathbf{P}^{C}_{ij}$ encodes the directional confidence from the LTI model, while $w_{ij}$ adaptively rescales the precision based on the observed residual $\mathbf{r}_{ij}$, down-weighting measurements with larger than expected deviations.
The metric $1/(1+\mathbf{r}^{\top}\mathbf{P}\mathbf{r}/\nu)$ behaves similarly to the unnormalized softmax: if one residual is much smaller than the others, it dominates, while if all residuals are similar and not too large, the output approximates a uniform weighting. The key difference arises when all residuals are small relative to $\nu$: unlike softmax, which picks a dominant term whenever relative differences are large, this similarity score reflects uncertainty and maintains a roughly uniform attention distribution. In this sense, $\nu$ acts as a kind of noise floor.
Using the fact that $\log(1+x)\approx x,\,\,\text{for }x\ll 1$, we can expand the rational form of $w_{ij}$:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}\approx\exp(-\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu).$
@@ Appendix B Additional Material
B.1 Separable Approximation of the Precision Tensor
@@
$\mathbf{D}[i,j]=\mathcal{P}^{C}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\mathcal{P}^{C}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}$
$-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathcal{P}^{C}[k,i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
Our previous approach — imposing a model with isotropic decay $\alpha$ and noise $\sigma,\eta$ — achieved $\mathcal{O}(m^{2})$ memory, thereby allowing the precision tensor $\mathcal{P}^{C}$ to be factored out of the summation over $k$. Here, we attempt to maintain the full $k$-dependence of the time-dependent precision tensor $\mathcal{P}^{C}[k,i,j]$ while still achieving $\mathcal{O}(m^{2}+md)$ memory. This is achieved by seeking a low-rank (separable) approximation of the precision kernel.
@@
$\mathcal{P}^{C}[k,i,j]\approx\mathbf{c}[k]+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j].$
@@
Recall that the precision tensor is defined by:
$\mathcal{P}^{C}[k,i,j]=\left(\lambda_{C,k}^{2}\odot\lambda_{Q,k}\odot\frac{1-e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}}{-2\mathrm{Re}(\lambda_{k})}+\lambda_{R,k}\odot e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}\right)^{-1}$
@@
$\mathcal{P}^{C}[k,i,j]=\left(\lambda_{\text{eff},k}+\big(-\lambda_{\text{eff},k}+\lambda_{R,k}\big)\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1}$
$\frac{1}{\lambda_{\text{eff},k}}\left(1+\lambda_{r,k}\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1},\quad\text{where }\lambda_{r,k}=-1+\frac{\lambda_{R,k}}{\lambda_{\text{eff},k}}>-1$
@@
$\mathcal{P}^{C}_{k}(\Delta t)\approx\frac{1}{\lambda_{\text{eff},k}}\left(1-\frac{\lambda_{r,k}}{1+\lambda_{r,k}}\,e^{\tfrac{2\,\mathrm{Re}(\lambda_{k})\,\Delta t}{1+\lambda_{r,k}}}\right)=\frac{1}{\lambda_{\text{eff},k}}\left[1-\left(1-\frac{\lambda_{\text{eff},k}}{\lambda_{R,k}}\right)e^{-\tfrac{|\lambda_{C,k}|^{2}\,\lambda_{Q,k}}{\lambda_{R,k}}\,\Delta t}\right]$
@@
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\frac{\mathcal{P}^{C}[k,i,j]}{1+\mathbf{D}[i,j]}\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
@@
While this formulation reduces the memory complexity from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, the reliance on numerous element-wise operations introduces a high constant factor to the $\mathcal{O}(m^{2}d)$ time complexity, making it less scalable than the isotropic decay model.
B.2 Recurrent Formulation
We now describe a way to incorporate state-dependent weighting from the full batch MLE while retaining the recursive structure of the Kalman Filter. The key idea is to use the innovation residual at each step to reweight the update, and iterate within each time step to approximate the effect of all pairwise comparisons. This iteration again takes place in the eigenbasis.
@@ B.2 Recurrent Formulation
While the batch MLE uses the global information provided by pair-wise residuals, the innovation-based weight $w_{i}$ in the recursive update depends only on the current residual $r_{i}=\mathbf{z}_{i}-\mathbf{\bar{z}}_{i}^{-}$. However, when the dynamics are slow (small eigenvalues), $\Delta t$ is sufficiently small, and the residuals are not too large, $\mathbf{\bar{z}}_{\mathbf{s}i}^{-}$ effectively encodes past measurements. In this case, with enough inner iterations, IRKF should approximate the batch IRLS solution, while retaining the recursive structure of the Kalman Filter. The number of pair-wise comparisons for a sequence of length $m$ is then reduced from $\mathcal{O}(m^{2})$ to $\mathcal{O}(Km)$, where $K\ll m$ is the number of inner iterations.
At inference time, we can approximately unroll our learned model into a reweighted Kalman Filter. The outer iteration propagates latent states through the sequence for fast, online inference, while the inner iterations at each step approximate the effect of full pairwise attention without incurring $\mathcal{O}(m^{2})$ computations. The recursive precision update acts as an implicit normalization of the "attention" weights.
@@
Lift real inputs to complex by concatenating zeros: $\mathbf{z}=\left[\mathbf{z}_{r},\mathbf{0}\right]^{\top}$
$\big(\operatorname{Re}(\mathbf{z_{v}}),\operatorname{Im}(\mathbf{z_{v}})\big)\leftarrow\mathcal{L}_{V}\big(\operatorname{Re}(\mathbf{z}),\operatorname{Im}(\mathbf{z})\big).$
@@
Innovation covariance: $\sigma^{2}_{V}=\sigma^{2}\kappa+\eta^{2}\Phi^{2}_{\alpha},\quad\kappa=\frac{1-\Phi^{2}_{\alpha}+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\Delta t\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
@@
    "Attention" weight: $w_{i}^{(k)}=-\log\left(1+p_{i}^{+(k)}\mathbf{y}_{\mathbf{v}i}^{(k)\dagger}\mathbf{y}_{\mathbf{v}i}^{(k)}+\epsilon\right)$
@@
Map back to original basis: $\big(\operatorname{Re}(\mathbf{z_{p}}),\operatorname{Im}(\mathbf{z_{p}})\big)\leftarrow\mathcal{L}_{P}\big(\operatorname{Re}(\mathbf{z}_{\mathbf{v}i}^{+}),\operatorname{Im}(\mathbf{z}_{\mathbf{v}i}^{+})\big)$
@@
B.3 Generative Modeling
B.3.1 Propagating the Covariance through the Prediction Step
In particular, given our current estimate for the $i$th measurement, $\mathbf{\bar{z}}_{i}(t_{j})$, the one-step forward prediction is:
$\mathbf{z}_{i}^{+}(t_{j}+\Delta t)=\mathbf{C}e^{\mathbf{A}\Delta t}\mathbf{C}^{-1}\mathbf{\bar{z}}_{i}(t_{j})$
In the eigenbasis, this simplifies to:
Consider the innovation of the prediction at the next time step: $\mathbf{y}_{\mathbf{s}i}(t_{j}+\Delta t)=\mathbf{\bar{z}}_{\mathbf{s}i}(t_{j}+\Delta t)-\mathbf{z}^{C}_{\mathbf{s}i}(t_{j}+\Delta t)$, i.e. the difference between our predicted measurement and the true measurement. We want to compute $\text{Cov}\big(\mathbf{y}_{\mathbf{s}i}(t_{j}+\Delta t)\big)$.
The innovation covariance at the next step has two contributions:
  1. 1.
  2. 2.
Suppose the current innovation covariance for the $i$th estimate is:
$\mathrm{Cov}\big(\mathbf{y}_{\mathbf{s}i}\big)=\bm{\Lambda}_{\mathrm{innov},i}^{-}.$
In particular,
$\mathrm{Cov}\big(\mathbf{y}_{\mathbf{s}i}\big)=\left(\sum_{j\leq i}w_{ij}\mathbf{P}^{C}_{ij}\right)^{-1},\quad\text{where:}\quad w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}.$
(We already computed this during the estimation step.)
Propagating this through the deterministic dynamics gives:
To sample a temporally consistent trajectory, we can use a form of backward recursive sampling based on the Rauch-Tung-Striebel (RTS) smoother: Starting at the final time step $i=m$ and moving backward to $i=1$, we sample the state: $\tilde{\mathbf{z}}_{i}$ conditioned on the previously sampled state $\tilde{\mathbf{z}}_{i+1}$ and the filtered estimate $\mathbf{z}_{i|i}$.
B.4 Model Predictive Control
B.4.1 Open loop waypoint control

### v3

@@ 7 Discussion and Conclusion
The core contribution of Adaptive Filter Attention (AFA) is the principled unification of attention mechanisms with adaptive filtering, and the inclusion of temporal structure and explicit uncertainty modeling. Whereas dot-product attention heuristically defines its weights, AFA derives them from parallel maximum likelihood estimates for each point in the trajectory, under a shared linear time-invariant stochastic differential equation (SDE) model. This framework yields a rigorous probabilistic interpretation of attention as precision-weighted averaging — essentially a parallelized form of the adaptive Kalman filter.
Using the assumption of simultaneously-diagonalizable linear dynamics and noise matrices, we obtain a tensor generalization of the attention mechanism. Imposing an isotropic-plus-skew-symmetric state matrix and isotropic noise, our model achieves the same asymptotic time and memory complexity as standard attention. Under the assumption of unitary-dynamics and zero process noise, AFA reduces to complex-valued dot-product attention with a rotary positional encoding, identifying the latter as a noise-ignorant special case of a broader, probabilistically grounded adaptive filtering framework.
Practically, AFA introduces a number of architectural changes that improve the expressivity and inductive biases of standard attention. It employs fully complex-valued computations to enable a generalized positional encoding that extends RoPE from fixed positional rotations to transformations governed by learned linear dynamics. It also performs rotation–unrotation of value vectors to yield dynamically consistent, position-aware state estimates rather than simple weighted averages. AFA’s uncertainty-aware weighting implements a robust optimal filter, while the continuous-time formulation naturally accommodates unequal time steps. Finally, AFA enforces dynamical self-consistency by using the same learned dynamics for both state/uncertainty propagation, and for estimation/prediction, providing a form of temporal regularization.
@@
We also introduce a Radial–Tangential SDE (RT-SDE) model, which generalizes Isotropic AFA by assuming the process and measurement noise decompose into separate radial and tangential components, in the token’s rotating frame. The RT-SDE provides an optimal trade-off between expressive power (allowing anisotropic noise) and computational efficiency, as it remains analytically tractable and fully scalable. This formulation leads directly to a maximum likelihood estimator on the hypersphere, which provides a novel, principled foundation for understanding norm-preserving attention mechanisms, and the architectural properties of the Transformer.
@@
Several promising directions remain for future work. A natural extension is to allow the dynamcis or control inputs to depend on the state. Although a fully time-varying dynamics model would break the linear time-invariant (LTI) structure and would no longer allow a closed-form solution to the differential Lyapunov equation, this problem can be avoided by fixing the dynamics within each context window. In this way, the parameters may adapt across windows while remaining constant within a window, enabling content adaptivity without sacrificing computational efficiency.
A natural extension is to try to relax the assumption of diagonalizable state matrix $\mathbf{A}$ to diagonalizable-plus-rank-1 (DPR1). If one assumes that the rank-1 component must be confined to a subspace where all the dimensions share a single rotational frequency, the propagated covariance maintains this DPR1 form, significantly simplifying the uncertainty dynamics. This structure enables the Mahalanobis distance to be efficiently calculated as a linear combination of the full residual norm and the residual projected onto the rank-1 subspace, and factorizes the optimal estimator into two parallel weighted-average computations, one for the uncoupled features and one for the dynamically coupled features, so that the entire mechanism should have the same asymptotic complexity as the diagonalizable case, but with a higher constant factor. We could also consider generalizations using the Jordan canonical form, differentiable numerical solvers for the differential Lyapunov equation, or unrolled iterations of specialized low-rank methods such as Krylov subspace techniques or ADI iterations.
@@
The duality between attention weights and the Kalman gain suggests a path toward dynamics-aware neural control policies, where the separation principle can be used to achieve more sample-efficient policy evaluation and state estimation in noisy or partially observed reinforcement learning settings.
Finally, AFA provides a natural interpretation of attention weights as pairwise precisions between queries and keys. Since these weights are derived from uncertainty propagation, they are a direct, statistically grounded measure of information reliability. Future work could investigate using this explicit measure of precision to improve model robustness or to visualize how uncertainty propagates through the network.
@@ Appendix A Background
A.1 Propagation of Uncertainty through a Linear SDE
@@ A.3 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\sigma_{V}^{2}(\Delta t_{ij})\mathbf{I},$
$\sigma_{V}^{2}(\Delta t_{ij}):=\text{sign}(\Delta t_{ij})\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}+e^{-2\alpha\Delta t_{ij}}\eta^{2}$
@@ A.4 Variational Formulation of Robust Estimation
Dividing numerator and denominator by $\beta$ and setting the numerator to $1$ (since it is canceled by the normalization), there is only one independent constant, $\nu=2\beta$, which controls the residual scaling, so the robust weight takes the form:
@@ Appendix B Additional Material
B.1 Generative Modeling
In particular, given our current estimate for the $i$th measurement, $\mathbf{\bar{z}}_{i}(t_{j})$, the one-step forward prediction in the eigenbasis is:
Consider the innovation of the prediction at the next time step: $\mathbf{y}_{\mathbf{s}i}(t_{j}+\Delta t):=\mathbf{\bar{z}}_{\mathbf{s}i}(t_{j}+\Delta t)-\mathbf{z}^{C}_{\mathbf{s}i}(t_{j}+\Delta t)$, i.e. the difference between our predicted measurement and the true measurement. We want to compute $\text{Cov}\big(\mathbf{y}_{\mathbf{s}i}(t_{j}+\Delta t)\big)$.
The innovation covariance at the next step has two contributions:
  1. 1.
  2. 2.
The current innovation covariance for the $i$th estimate is:
$\bm{\Lambda}_{\mathrm{innov},i}^{-}:=\mathrm{Cov}\big(\mathbf{y}_{\mathbf{s}i}\big)=\left(\sum_{j\leq i}w_{ij}\mathbf{P}^{C}_{ij}\right)^{-1}=\left(\sum_{j\leq i}\mathbf{\tilde{A}}[i,j]\right)^{-1}\mathbf{I}$
(We already computed this during the estimation step.)
Propagating this through the deterministic dynamics gives:
In the case of isotropic decay and noise, this is simply:
$\bm{\Lambda}_{i}^{+}=e^{2\alpha\Delta t}\,\bm{\Lambda}_{\mathrm{innov},i}^{-}+\|\lambda_{C}\|^{2}\sigma^{2}\frac{1-e^{2\alpha\Delta t}}{-2\alpha}+\eta^{2}$
While the above method is suitable for sequential, causal next-token generation, it only defines the probability of the next step. If a complete, temporally consistent past trajectory is required, a backward recursive sampling scheme based on the Rauch-Tung-Striebel (RTS) smoother can be employed. The RTS procedure starts at the final time step $i=m$ and moves backward to $i=1$, sampling the state $\mathbf{\tilde{z}}_{\mathbf{s},i}$ conditioned on the previously sampled state $\mathbf{\tilde{z}}_{\mathbf{s},i+1}$ and the forward-filtered estimate $\mathbf{z}_{\mathbf{s},i|i}$.
$\mathbf{\tilde{z}}_{\mathbf{s},i}\sim p(\mathbf{z}_{\mathbf{s},i}\,|\,\mathbf{\tilde{z}}_{\mathbf{s},i+1},\mathbf{z}_{\mathbf{s},i|i})$
Note, however, that this backward pass is computationally intensive (typically $\mathcal{O}(md^{2})$ per layer).
B.2 Confidence-Aware Gating
AFA provides a probabilistic interpretation of the attention weights: they represent the pair-wise precisions predicted by the dynamics model, reweighted by a robust weight based on the pair-wise residuals. Hence, the row-sum of the unnormalized attention matrix $\sum_{j}\mathbf{\tilde{A}}[i,j]$ is a measure of the total confidence of a particular query.
By construction, the normalization step in AFA, as in other attention mechanisms, enforces the total attention mass to equal one ($\sum_{j}\mathbf{A}[i,j]=1$), effectively removing a query’s ability to communicate its overall confidence. As a result, the model cannot distinguish between two distinct uncertainty regimes: high-confidence uncertainty, where many keys provide equally strong evidence and averaging yields a precise consensus, and low-confidence uncertainty, where all keys are poor matches and the average merely reflects noise.
To address this limitation, we can extract the row-sums of the unnormalized attention score (i.e. the information matrix) and transform this global evidence into a feature-wise confidence gate that scales the final attention output. This mechanism enables each value to modulate its contribution to the residual stream in proportion to the total statistical support it receives.
We define a gate as a scalar confidence factor for value $i$, using a gating function $\sigma$:
$\bm{\Gamma}[i]=\sigma\!\left(w_{\Gamma}\sum_{j}\mathbf{\tilde{A}}[i,j]+b_{\Gamma}\right)$
where $\mathbf{\tilde{A}}$ is the unnormalized attention matrix and $w_{\Gamma},b_{\Gamma}\in\mathbb{R}$ are learned scalar weights.
The confidence gate modulates the final estimated state:
$\mathbf{\bar{Z}}_{\text{output}}[i,k]=(1-\delta)\,\bm{\Gamma}[i]\cdot\mathbf{\bar{Z}_{v}}[i,k]+\delta\,\mathbf{Z_{v}}[i,k]$
This operation can be computed efficiently because it reuses the row-sum of the unnormalized attention matrix $\mathbf{\tilde{A}}$, which has already been calculated.
In the IRLS interpretation, the confidence-aware gate $\bm{\Gamma}$ scales the update, allowing the model to more aggressively update tokens with high-precision estimates and dampen updates for less certain ones, which stabilizes the iteration. This policy is crucial because the input $\mathbf{Z}_{\mathbf{v}}$ represents the robustly filtered state from all preceding layers, while $\mathbf{\bar{Z}}_{\mathbf{v}}$ is only the single-step estimate from the current layer’s filter. Hence, when the total statistical evidence is weak, the update is suppressed, and the output conservatively reverts to the proven, previously refined state.
B.3 Separable Approximation of the Precision Tensor
@@
$\mathbf{D}[i,j]=\mathcal{P}^{Z}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\mathcal{P}^{Z}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}$
$-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathcal{P}^{Z}[k,i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
Our previous approach — imposing a model with isotropic decay $\alpha$ and noise $\sigma,\eta$ — achieved $\mathcal{O}(m^{2})$ memory, thereby allowing the precision tensor $\mathcal{P}^{Z}$ to be factored out of the summation over $k$. Here, we attempt to maintain the full $k$-dependence of the time-dependent precision tensor $\mathcal{P}^{Z}[k,i,j]$ while still achieving $\mathcal{O}(m^{2}+md)$ memory. This is achieved by seeking a low-rank (separable) approximation of the precision kernel.
@@
$\mathcal{P}^{Z}[k,i,j]\approx\mathbf{c}[k]+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j].$
@@
Recall that the precision tensor is defined as:
$\mathcal{P}^{Z}[k,i,j]=\left(\lambda_{C,k}^{2}\lambda_{Q,k}\frac{1-e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}}{-2\mathrm{Re}(\lambda_{k})}+\lambda_{Rk,k}e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}+\lambda_{Rq,k}\right)^{-1}$
@@
$\mathcal{P}^{Z}[k,i,j]=\left(\lambda_{\text{eff},k}+\lambda_{Rq,k}+\big(-\lambda_{\text{eff},k}+\lambda_{Rk,k}\big)\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1}$
$\frac{1}{\lambda_{\text{eff},k}+\lambda_{Rq,k}}\left(1+\lambda_{r,k}\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1},\quad\text{where }\lambda_{r,k}=\frac{-\lambda_{\text{eff},k}+\lambda_{Rk,k}}{\lambda_{\text{eff},k}+\lambda_{Rq,k}}>-1$
@@
$\mathcal{P}^{Z}_{k}(\Delta t)\approx\frac{1}{\lambda_{\text{eff},k}+\lambda_{Rq,k}}\left(1-\frac{\lambda_{r,k}}{1+\lambda_{r,k}}\,e^{\tfrac{2\,\mathrm{Re}(\lambda_{k})\,\Delta t}{1+\lambda_{r,k}}}\right)$
@@
$\mathbf{\bar{Z}_{v}}[k,i]=\bm{\Phi}_{\mathbf{v}}[k,i]\cdot\sum_{j\leq i}\frac{\mathcal{P}^{Z}[k,i,j]}{1+\mathbf{D}[i,j]}\cdot\mathbf{\hat{Z}}_{\mathbf{v}}[k,j]$
@@
While this formulation reduces the memory complexity from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, the reliance on numerous element-wise operations introduces a high constant factor to the $\mathcal{O}(m^{2}d)$ time complexity, making it less scalable than the isotropic decay model. Moreover, it is not self-consistent, since a bias is introduced into the precision, which may interfere with the proper learning of the underlying dynamics model.
B.4 Recurrent Formulation
We now describe a way to incorporate state-dependent weighting from the full batch MLE while retaining the recursive structure of the Kalman Filter. The key idea is to use the innovation residual at each step to reweight the update (in the eigenbasis), and iterate within each time step to approximate the effect of all pairwise comparisons.
@@ B.4 Recurrent Formulation
While the batch MLE uses the global information provided by pair-wise residuals, the innovation-based weight $w_{i}$ in the recursive update depends only on the current residual $r_{i}=\mathbf{z}_{i}-\mathbf{\bar{z}}_{i}^{-}$. However, when the dynamics are slow (small eigenvalues), $\Delta t$ is sufficiently small, and the residuals are not too large, $\mathbf{\bar{z}}_{\mathbf{s}i}^{-}$ effectively encodes past measurements. In this case, with enough inner iterations, IRKF should approximate the batch IRLS solution, while retaining the recursive structure of the Kalman Filter. The number of pair-wise comparisons for a sequence of length $m$ is then reduced from $\mathcal{O}(m^{2})$ to $\mathcal{O}(Km)$, where $K\ll m$ is the number of inner iterations. This effectively strikes a middle ground between the full attention mechanism and a state space model, as it uses a recurrent structure, while retaining non-linear, state-dependent weighting.
@@
$\big(\operatorname{Re}(\mathbf{z_{v}}),\operatorname{Im}(\mathbf{z_{v}})\big)\leftarrow\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{z}\big).$
@@
Innovation covariance: $\sigma^{2}_{V}=\sigma^{2}\kappa+\eta^{2}\Phi^{2}_{\alpha}+\gamma^{2},\quad\kappa=\frac{1-\Phi^{2}_{\alpha}+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\Delta t\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
@@
    "Attention" weight: $w_{i}^{(k)}=\left(1+p_{i}^{+(k)}\mathbf{y}_{\mathbf{v}i}^{(k)\dagger}\mathbf{y}_{\mathbf{v}i}^{(k)}+\epsilon\right)^{-1}$
@@
Output projection:
$\operatorname{Re}(\mathbf{z_{p}})\leftarrow\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}\big(\operatorname{Re}(\mathbf{z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{z}_{\mathbf{v}}^{+})\big)$
@@
B.5 Model Predictive Control